A distribution of values is normal with a mean of 231.6 and a standard deviation of 62.4.
Find the probability that a randomly selected value is less than 312.7.
P(X < 312.7) =

Answer by stanbon(75887)

A distribution of values is normal with a mean of 231.6 and a standard deviation of 62.4.
Find the probability that a randomly selected value is less than 312.7.
P(X < 312.7) =
----
z(312.7) = (312.7-231.6)/62.4 = 1.3000
----
P(x < 312.7) = P(z < 1.3000) = 0.9031
================
Cheers,
Stan H.
================